The current I in an electrical conductor varies inversely as the resistance R of the conductor. The current is 1/7 ampere when the resistance is 35015 ohms. What is the current when the resistance is 18375 ohms?
Answer by stanbon(48569)

The current I in an electrical conductor varies inversely as the resistance R of the conductor.
I = k/R
----
Find k if the current is 1/7 ampere when the resistance is 35015 ohms.
(1/7) = k/35015
k = 5002.14
---
So equation is I = 5002.14/R
---
What is the current when the resistance is 18375 ohms?
I = 5002.14/18375
I = 0.27222...ampere
============================
Cheers,
Stan H.
---------------------------------